# Consumption Partial Insurance

### v5

@@ B.4.3 Moments of consumption – quadratic function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2(\phi^{(2)})^{2}(\kappa_{\zeta}+3(\sigma_{\zeta}^{2})^{2})+(\psi^{(2)})^{2}\kappa_{v}$
$+2(\omega^{(22)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}+4\phi^{(1)}\phi^{(2)}\gamma_{\zeta}+2\psi^{(1)}\psi^{(2)}\gamma_{v}+4\phi^{(2)}\psi^{(2)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ $=4(\phi^{(2)})^{2}(\sigma_{\zeta}^{2})^{2}+4\phi^{(2)}\psi^{(2)}\sigma_{\zeta}^{2}\sigma^{2}_{v}+(\psi^{(2)})^{2}(\sigma^{2}_{v})^{2}-\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}+2\phi^{(2)}\gamma_{\zeta}+\psi^{(2)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\sigma_{v}^{2}-\psi^{(2)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\gamma_{\zeta}+\psi^{(1)}\gamma_{v}+2\phi^{(2)}(\kappa_{\zeta}+3(\sigma_{\zeta}^{2})^{2}+2\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2})$
$+\psi^{(2)}(\kappa_{v}+(\sigma_{v}^{2})^{2}+2\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{v}^{2}\sigma_{u^{y}}^{2})+4\omega^{(22)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\gamma_{v}+4\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})+\psi^{(2)}(\kappa_{v}+\sigma_{v}^{2}(2\sigma_{\zeta}^{2}+\sigma_{v}^{2}+2\sigma_{u^{y}}^{2}))$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=4\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})+2\psi^{(2)}\sigma_{v}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\gamma_{v}-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}(\kappa_{v}+\sigma_{v}^{2}\sigma_{u^{y}}^{2})-2\omega^{(22)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ $=-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ $=-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
@@ C.1 Summary statistics in the PSID sample
Notes: The table reports summary statistics for consumption, income, wealth, and demographics in our baseline sample in the PSID and in subsamples sorted by wealth and education, over 19991999199919992019201920192019. Monetary amounts are expressed in \$2018. Rent for owners is calculated at 6% of the self-reported value of the house.
@@ C.2 Comparison with administrative moments
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.065 (0.004)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.569 (0.128) -1.039
@@
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.056 (0.004)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.728 (0.142)
@@ C.4 CEX data and imputation in the PSID
Suppose that consumption and income measurement errors are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{c},u_{it}^{y})=\sigma_{u_{t}}>0$. The positive correlation may be the result of, e.g., recall bias that affects similarly all monetary amounts. This is something we have not allowed for in the paper, and neither have BPP. While $\phi_{t}^{(1)}$ is identified via ${\operatorname*{\mathbb{E}}}(\Delta c_{it}(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, in practice BPP and virtually all follow-up papers estimate it through $({\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it})+{\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, assuming that $\sigma_{u_{t}}=0$ (uncorrelated errors). This is what we also do in our baseline estimation of $\phi_{t}^{(1)}$ in the linear specification. However, if the errors are correlated, then this target moment in fact identifies $\phi_{t}^{(1)\text{psid*}}=\phi_{t}^{(1)}+\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. Clearly, correlation between consumption and income errors causes positive bias as $\phi_{t}^{(1)\text{psid*}}>\phi_{t}^{(1)}$. This is true in the baseline, i.e., without the use of imputed consumption, but the bias becomes multiplicatively larger with imputed consumption, as we show below.
@@
Suppose that food expenditure in the PSID is measured with error, so imputed consumption is $c_{it}^{\text{imp*}}=\kappa_{t}(f_{it}+u_{it}^{f})$. Suppose that food consumption and income error are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{f},u_{it}^{y})=\sigma_{u_{t}}>0$. Because $\kappa_{t}>1$, the consumption imputation magnifies the role of the consumption measurement error. One can show that, with imputed consumption, the moment used to identify $\phi_{t}^{(1)}$ in fact identifies $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
@@ C.5 Estimation using the original BPP data
#$\theta$ is the moving average parameter of the transitory shock. BPP assume that the transitory component of income is given by $v_{it}=\epsilon_{it}+\theta\epsilon_{it-1}$, where $\epsilon_{it}$ is the transitory shock. Given their use of annual growth rates, $\theta$ is identified by the first-order income autocovariance. Our use of biennial data in this paper rules out identification of $\theta$, which we thus set to 0.
@@ C.6 Additional results
Tables C.6 and C.7 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix in the GMM estimation (optimal GMM). As such, these tables accompany tables 2 and 3 in the paper, which report results from equally weighted GMM. The results are qualitatively (and most often numerically) similar to those reported in the paper.
@@
Table C.8 presents the estimates of the parameters of the consumption function using imputed consumption data from the CEX. As such, this table accompanies table 3 in the main text, which reports results using consumption internally available in the PSID. In the linear specification, the transmission parameter of permanent shocks $\phi^{(1)}$ is consistently higher than that obtained from the consumption data internally available in the PSID, regardless of whether we target only second- or also higher-order moments. This echoes our earlier discussion about the imputation inflating the pass-through of shocks. We also find that the transmission parameter of transitory shocks $\psi^{(1)}$ is negative in all cases. Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude and negative ($-0.418$). This is counterintuitive, raising further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved. We draw similar conclusions about $\phi^{(1)}$ and $\psi^{(1)}$ in the context of the quadratic function. $\phi^{(2)}<0$ in that case, with a magnitude similar to our baseline results from the PSID ($-0.036$ versus $-0.04$; albeit not significant). Once again, this is indicative of the asymmetric pass-through of negative versus positive permanent shocks. Similar to our baseline, we also find $\psi^{(2)}>0$.
@@
Table C.9 presents the estimates of the parameters of the income process by age and the household’s position in the income distribution. Skewness of all shocks becomes more negative, both economically and statistically, with age. Skewness of permanent shocks is weakly positive in the bottom income tertile but turns very negative in the top tertile. The opposite is true for transitory shocks; skewness is negative in the bottom income tertile but strongly positive in the top.
@@
Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution, respectively. Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65. A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted. This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results. [^7] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters. By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details. By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets. Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, have more insurance against permanent income shocks than the poor.
@@
Table C.12 presents estimation results for the linear consumption function in the subsamples formed by wealth and education. As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification. The odd-numbered columns present parameter estimates when only second-order moments are targeted. $\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across groups. This is similar to what we found in the quadratic specification. However, as soon as we target up to fourth-order moments in the linear specification, shown in the even-numbered columns, this monotonic pattern is lost. Transitory shocks transmit strongly in group 1, and with a negative coefficient.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary by wealth and education. The underlying income moments also vary by wealth and education (results not shown for brevity). We do not report results when up to third-order moments of income and consumption are targeted, but these are available upon request. Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No college versus some college is defined on the basis of the highest level of education attained by the male spouse (cutoff is 12 years of schooling). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. This table accompanies the quadratic specification estimates in table 6 in the paper.

### v6

@@ B.4.3 Moments of consumption – quadratic function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+(\phi^{(2)})^{2}\Big((1+\rho^{4})\kappa_{\zeta}+6\rho^{2}(\sigma_{\zeta}^{2})^{2}\Big)$
$+(\psi^{(2)})^{2}\kappa_{v}+(\omega^{(22)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}$
$+2\phi^{(1)}\phi^{(2)}(1+\rho^{3})\gamma_{\zeta}+2\psi^{(1)}\psi^{(2)}\gamma_{v}+2\phi^{(2)}\psi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ $=(\phi^{(2)})^{2}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+2\phi^{(2)}\psi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma^{2}_{v}+(\psi^{(2)})^{2}(\sigma^{2}_{v})^{2}-\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}+\phi^{(2)}(1+\rho^{3})\gamma_{\zeta}+\psi^{(2)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\rho^{2}\sigma_{v}^{2}-\psi^{(2)}\rho^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{3})\gamma_{\zeta}+\psi^{(1)}\gamma_{v}$
$+\phi^{(2)}\Big((1+\rho^{4})\kappa_{\zeta}+6\rho^{2}(\sigma_{\zeta}^{2})^{2}+(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
$+\psi^{(2)}\Big(\kappa_{v}+\rho^{4}(\sigma_{v}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}\Big)+2\omega^{(22)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\rho^{4}\gamma_{v}+\phi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\Big((1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
$+\psi^{(2)}\Big(\rho^{4}\kappa_{v}+(\sigma_{v}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}\Big)$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=\phi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\Big((1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
$+\psi^{(2)}\sigma_{v}^{2}\Big((1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\rho^{2}\gamma_{v}-\phi^{(2)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
$-\psi^{(2)}\rho^{2}(\kappa_{v}+\sigma_{v}^{2}\sigma_{u^{y}}^{2})-\omega^{(22)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ $=-\phi^{(2)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\rho^{2}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ $=-\phi^{(2)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\rho^{2}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
@@ C.1 Summary statistics in the PSID sample
Notes: The table reports summary statistics for consumption, income, wealth, and demographics in our baseline sample in the PSID and in subsamples sorted by wealth and education, over 1999199920192019. Monetary amounts are expressed in \$2018. Rent for owners is calculated at 6% of the self-reported value of the house.
@@ C.2 Comparison with administrative moments
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$ -0.065 (0.004)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$ -0.569 (0.128) -1.039
@@
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$ -0.056 (0.004)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$ -0.728 (0.142)
@@ C.4 CEX data and imputation in the PSID
Suppose that consumption and income measurement errors are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{c},u_{it}^{y})=\sigma_{u_{t}}>0$. The positive correlation may be the result of, e.g., recall bias that affects similarly all monetary amounts. This is something we have not allowed for in the paper, and neither have BPP. While $\phi_{t}^{(1)}$ is identified via ${\operatorname*{\mathbb{E}}}(\Delta c_{it}(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$ in the baseline permanent-transitory income process, in practice BPP and virtually all follow-up papers estimate it through $({\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it})+{\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, assuming that $\sigma_{u_{t}}=0$ (uncorrelated errors). This is what we also do in our baseline estimation of $\phi_{t}^{(1)}$ in the linear specification. However, if the errors are correlated, then this target moment in fact identifies $\phi_{t}^{(1)\text{psid*}}=\phi_{t}^{(1)}+\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. Clearly, correlation between consumption and income errors causes positive bias as $\phi_{t}^{(1)\text{psid*}}>\phi_{t}^{(1)}$. This is true in the baseline, i.e., without the use of imputed consumption, but the bias becomes multiplicatively larger with imputed consumption, as we show below.
@@
Suppose that food expenditure in the PSID is measured with error, so imputed consumption is $c_{it}^{\text{imp*}}=\kappa_{t}(f_{it}+u_{it}^{f})$. Suppose that food consumption and income error are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{f},u_{it}^{y})=\sigma_{u_{t}}>0$. Because $\kappa_{t}>1$, the consumption imputation magnifies the role of the consumption measurement error. One can show that, with imputed consumption, the moment used to identify $\phi_{t}^{(1)}$ in fact identifies $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. As a result, $\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
@@ C.5 Estimation using the original BPP data
$\theta$ is the moving average parameter of the transitory shock. BPP assume that the transitory component of income is given by $v_{it}=\epsilon_{it}+\theta\epsilon_{it-1}$, where $\epsilon_{it}$ is the transitory shock. Given their use of annual growth rates, $\theta$ is identified by the first-order income autocovariance. Our use of biennial data in this paper rules out identification of $\theta$, which we thus set to 0.
@@ C.6 Additional results
Tables C.6 and C.7 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix (optimal GMM). As such, these tables accompany tables 2 and 3 in the paper, which report results from equally weighted GMM.
@@
Table C.8 presents the estimates of the parameters of the consumption function using imputed consumption from the CEX. As such, this table accompanies table 3 in the paper, which reports results using consumption internally available in the PSID. In the linear specification, $\phi^{(1)}$ is consistently higher than that obtained from the consumption data internally available in the PSID, regardless of whether we target only second- or also higher-order moments. This echoes our earlier discussion about the imputation inflating the pass-through of shocks. We also find that $\psi^{(1)}$ is negative in all cases. Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude (0.418-0.418). This is counterintuitive, raising further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved. We draw similar conclusions about $\phi^{(1)}$ and $\psi^{(1)}$ in the context of the quadratic function. $\phi^{(2)}<0$ in that case, with a magnitude similar to our baseline results from the PSID (0.036-0.036 versus 0.04-0.04; albeit not significant). Once again, this is indicative of the asymmetric pass-through of negative versus positive permanent shocks. Similar to our baseline, we find $\psi^{(2)}>0$.
@@
Table C.9 presents the estimates of the income process by age and the household’s position in the income distribution. Skewness of all shocks is more negative at older ages. Skewness of permanent shocks is weakly positive in the bottom income tertile but turns very negative in the top tertile. The opposite is true for transitory shocks; skewness is negative in the bottom income tertile but strongly positive in the top.
@@
Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution, respectively. $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65. A similar non-monotonic pattern appears in column 2, where third moments are also targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted. This is indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results. [^8] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters. By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details. By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets. $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, have more insurance against permanent income shocks than the poor.
@@
Table C.12 presents estimation results for the linear consumption function in the subsamples formed by wealth and education. The odd-numbered columns present parameter estimates when only second-order moments are targeted. $\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across groups. This is similar to what we find in the quadratic specification. However, as soon as we target up to fourth-order moments in the even-numbered columns, this monotonic pattern is lost. Transitory shocks transmit strongly in group 1, and with a negative coefficient.
Table C.13 presents the estimates of the income process, when permanent income follows an AR(1) process, i.e., it features linear persistence. Table C.14 presents the associated estimates of the parameters of the linear and quadratic consumption functions.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary by wealth and education. The underlying income moments also vary by wealth and education (results not shown for brevity). We do not report results when up to third-order moments of income and consumption are targeted, but these are available upon request. Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 1999199920192019. No college versus some college is defined on the basis of the highest level of education attained by the male spouse (cutoff is 12 years of schooling). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. This table accompanies the quadratic specification estimates in table 6 in the paper.
Table C.13: Estimates of the income process, income persistence
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
Income data: PSID PSID PSID
(1) (2) (3)
Panel A. Permanent shocks
Variance $\sigma_{\zeta}^{2}$ 0.058 (0.016) 0.056 (0.016) 0.036 (0.006)
Skewness $\gamma_{\zeta}$ central -0.007 (0.003) -0.006 (0.003)
standardized -0.503 (0.337) -0.910 (0.512)
Kurtosis $\kappa_{\zeta}$ central 0.039 (0.007)
standardized 29.891 (11.459)
Panel B. Transitory shocks
Variance $\sigma_{v}^{2}$ 0.011 (0.015) 0.014 (0.014) 0.030 (0.005)
Skewness $\gamma_{v}$ central -0.005 (0.003) -0.006 (0.003)
standardized -3.272 (5.415) -1.178 (0.694)
Kurtosis $\kappa_{v}$ central 0.047 (0.006)
standardized 52.091 (19.426)
Panel C. Linear persistence
AR(1) $\rho$ 0.846 (0.038) 0.854 (0.038) 0.931 (0.028)
Notes: The table presents the estimates of the parameters of the AR(1) income process. Column 1 targets only the second moments of income growth; column 2 targets also third moments; column 3 targets all moments up to fourth-order. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. This table accompanies table 7 in the paper, as well as table C.14 that follows.
Table C.14: Estimates of the consumption function, income persistence
Consumption fn.: Linear Quadratic
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID PSID PSID PSID
(1) (2) (3) (4)
$\phi^{(1)}$ 0.089 0.090 0.110 0.112
(0.029) (0.029) (0.041) (0.024)
$\psi^{(1)}$ -0.139 -0.077 -0.013 -0.010
(0.265) (0.083) (0.077) (0.066)
$\phi^{(2)}$ -0.049
$\psi^{(2)}$ 0.019
$\omega^{(22)}$ 0.430
$\sigma^{2}_{\xi}$ 0.019 0.019 0.020 0.019
(0.001) (0.001) (0.001) (0.001)
$\gamma_{\xi}$ -0.001 -0.001
(0.001) (0.001)
$\kappa_{\xi}$ 0.031
$\sigma^{2}_{u_{c}}$ 0.044 0.044 0.044 0.044
(0.002) (0.002) (0.001) (0.002)
$\gamma_{u_{c}}$ 0.002 0.002
(0.001) (0.001)
$\kappa_{u_{c}}$ 0.013
Notes: The table presents the estimates of the parameters of the consumption function, when income features linear persistence. Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. This table accompanies table 7 in the paper.